Swift Oil Company is considering investing in a new oil well. It is expected that the oil well will increase annual revenues by $134,000 and will increase annual expenses by $76,000 including depreciation. The oil well will cost $449,000 and will have a $11,000 salvage value at the end of its 10-year useful life. Calculate the annual rate of return. (Round answer to 2 decimal places, e.g. 12.47.)
Answer #1

Annual rate of return is have two formula one is based on average investment and other is on initial investment

Average investment = (449000+11000/2) = 230000

Net income = 134000-76000 = 58000

a) Annual rate of return = net income/Average investment = 58000/230000 = 25.22%

b) Annual rate of return = Net income/Initial investment = 58000/449000 = 12.92%